crispyrat:

if x+y is 13 and x*y is 24 then what is the distance from the point to the origin

snowflake0531:

So we have x+y=13 xy=24 We have to solve for x and y, can you solve that?

snowflake0531:

Or, I suggest graphing it

crispyrat:

hm how would one graph this

snowflake0531:

Yea my fault, youc an't xd

crispyrat:

:( then what do you reccomened doing? should u try to solve for x and y?

snowflake0531:

Use substitution

crispyrat:

ok so x=13-y and then y(13-y)=13y-y^2=24

snowflake0531:

So that would make y^2 -13y +24

crispyrat:

should i do the same for x

snowflake0531:

Yea

crispyrat:

so 13-x=y x(13-x)=13x-x^2 so x is x^2-13x+24

crispyrat:

then since they equal the same thing i compare the to sides to get x and plug in one to y then i find the distance

crispyrat:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @crispyrat so 13-x=y x(13-x)=13x-x^2 so x is x^2-13x+24 \(\color{#0cbb34}{\text{End of Quote}}\) wait im a bit confused cus if they were the same then x and y would be the same xp

snowflake0531:

I think i'm more confused than you-

crispyrat:

so i tried doing thesame thing with x and got x+y=13 and so 13-x=y x(13-x)=13x-x^2 so x is x^2-13x+24

snowflake0531:

yes yes, ik, that's why i'm confused at that too-

snowflake0531:

@darkknight

snowflake0531:

@Angle

Angle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @crispyrat so i tried doing thesame thing with x and got x+y=13 and so 13-x=y x(13-x)=13x-x^2 so x is x^2-13x+24 \(\color{#0cbb34}{\text{End of Quote}}\) From this step, you can use the quadratic formula to solve for the x value. Then use the x value to find the y value. \(x = \huge \frac{-b\pm\sqrt{b^2-4ac}}{2a}\)

Angle:

Here are some other resources you can take a look at: https://www.wolframalpha.com/input/?i=x%2By%3D13+%3B+xy%3D24 And Desmos to figure out where the two equations intersect:

1 attachment